# Entropy Collapse in Mobile Sensors: The Hidden Risks of Sensor-Based Security

### v9

@@ 1 Introduction
Modern mobile devices expose rich motion and ambient signals from accelerometers, gyroscopes, magnetometers, and more. Beyond activity recognition and context inference, e.g. [1, 2, 3, 4], a long line of work proposes using these signals for security-critical tasks. This includes cryptographic key generation, zero-interaction device pairing, and continuous authentication [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Traditional key generation protocols [8, 18] have historically relied on shared motion patterns in order to establish secure communication between co-located devices. Meanwhile, other proposals have incorporated environmental phenomena, such as magnetic field characteristics and thermal fluctuations, to mitigate relay attacks [9, 10, 21, 5, 11] and to reduce the frequency of user authentication prompts [17, 16, 12, 13, 14].
Despite these advances, a crucial assumption remains under-examined: that sensor data provides sufficient entropy to thwart adversarial inference. Many existing studies rely on heuristic assessments or model classification metrics, e.g. [21, 7, 6, 9, 22, 19, 20], which do not address two basic questions: (1) how much entropy do sensors actually provide and (2) to what extent do multimodal sensor combinations improve entropy? For the latter, multimodal sensing has been proposed as a means to counteract sensor-specific weaknesses [9, 21, 22, 7], but the quantitative security benefits of combining multiple sensors have not been rigorously evaluated.
@@
The security implications of low-entropy sensor data can be both direct and severe. If such data is used as authentication evidence or to derive secret keys, an adversary may attempt to enumerate the entire space of possible measurement values in order to (re-)derive it. For such reasons, modern security standards provide important quantitative thresholds for what can be considered ‘secure’. For instance, NIST SP 800-57 [23] states that security levels below 80 bits are considered insecure for long-term use, with 112 bits as the current minimum and 128 bits recommended for cryptographic keys. Standards that govern random number generators (RNGs) are stricter still. NIST SP 800-90B [24] effectively requires $\approx$1 bit of min-entropy per output bit ($\epsilon\leq 2^{-32}$), while AIS 20/31 [25] mandates over 240 and 250 bits of min-entropy and Shannon entropy respectively for an RNG’s internal state. As we will demonstrate, the min-entropy of even complex combinations of mobile sensors falls well below these thresholds.
@@
In this paper, we analyse the entropy of 25 mobile sensors, which is substantially more than prior work—[26] (1 sensor), [27] (10), [17] (2), [28] (3)—across multiple public datasets. Our results reveal structural limits. Commodity sensors exhibit pronounced biases and dependencies, yielding on average 3.408–4.483 bits of min-entropy and 5.584–9.266 bits of Shannon entropy per sample across datasets. Multimodal fusion confers only modest gains: inter-sensor correlations and statistical bias both contribute to a divergence between average- and worst-case entropy, with min-entropy typically $\approx$40–75% below Shannon entropy for the same signals. In practical terms, schemes that treat these measurements as secrets expose a guess space small enough to be enumerable by an adversary. Our contributions are as follows:
@@ 2.1 Related Work
Mobile sensing has been widely explored for security tasks in three broad families: (1) proximity detection and relay-attack mitigation, where co-presence is inferred from shared motion or ambient context [29, 7, 11, 5, 21, 6, 9]; (2) device pairing, where two devices derive evidence of co-location to establish a secure association or shared secret [8, 30, 18, 20]; and (3) continuous authentication, which aims to verify users passively during normal interaction [16, 12, 13, 14, 31, 32, 33, 34]. Representative designs include tap-to-pair schemes using accelerometers [7], induced shared vibrations [6], and ambient-signal approaches that leverage temperature, humidity, magnetic field, or other environmental cues [9, 21]. These mechanisms commonly fuse accelerometer and gyroscope measurements to capture correlated motion as evidence of co-presence [8, 30].
@@
Our key observation is that model-performance metrics are typically decoupled from the entropy of the underlying data sources upon which such models are based. Across the aforementioned application areas, effectiveness is usually inferred from discriminative performance. A common pipeline—documented across proximity detection, pairing, and continuous authentication—collects data from $N$ participants; derives time- or frequency-domain features (e.g. cross-correlation, spectral energy, and Hamming, Euclidean, and mean absolute distances); learns a decision rule (thresholds or supervised models such as SVMs, Random Forests, Naïve Bayes); and reports FPR/FNR, EER, ROC/AUC, accuracy, or F1. This template appears in numerous systems, including [7, 9, 21, 11, 16, 14, 13, 31, 32, 6]. For a comprehensive treatment of the model evaluation metrics used by existing mobile sensor-based security systems, the reader is referred to the surveys by Abuhamad et al. [19], Alzubaidi and Kalita [35], and Mayrhofer and Sigg [20]. While discriminative performance metrics are informative for usability and functionality, they do not quantify entropy of the underlying sources, which is the focus of our analysis.
Adjacent evidence also highlights practical limits under tight timing constraints. For NFC transactions, EMV specifies $\approx$500 ms for proximity checks. In that setting, Markantonakis et al. [21] (building on [5, 11]) report EERs of 0.179–0.246 across sensor combinations, concluding that commodity sensors are unsuitable for reliable proximity and relay detection without incurring usability or security issues. Taken together, the literature demonstrates functional feasibility but leaves open the core question of how unpredictable the underlying measurements really are, especially under multimodal fusion. This motivates our entropy-centric evaluation in the remainder of the paper.
Some proposals have attempted to assess the entropy of sensor signals within the context of a security mechanism, but this represents a small minority of the literature. Specifically, T2Pair by Li et al. [15] reports 32.3–38.5 bits of Shannon entropy, and a refinement by Wu et al. [36] reports 51–54 bits. Even in the best cases, these values are low relative to modern cryptographic standards; moreover, we are aware of no work using standard min-entropy as a worst-case measure of unpredictability [24, 25, 37] (see comparison tables of state-of-the-art schemes in [35, 18, 20]). We address this gap by analysing the entropy of a wide range of mobile sensors, both in isolation and in combination.
@@
Concurrent work has attempted to assess the entropy bestowed by mobile sensors as a separate line of inquiry. Voris et al. [26] investigated accelerometers as true random number generators (TRNGs) on a WISP RFID tag and Nokia N97 phone. The authors find that min-entropy is proportional to the motion applied to the device, with stationary movement having the lowest min-entropy. Intrinsic noise from the sensor’s circuitry and seismic activity, and the sampling rate of its analog-to-digital converter (ADC), are found to influence entropy generation. Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer. Lv et al. [28] analysed three mobile sensors—a triaxial accelerometer, gyroscope, and magnetometer—on an undisclosed Xiaomi Redmi smartphone. Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method. Krhovják et al. [17] examined image and audio data collected from mobile phone cameras and microphones respectively on a Nokia N73 and E-Ten X500 and M700 phones. The work reports Shannon entropies of 2.9 (microphone) and 2.408–5.376 (camera), and min-entropies of 0.5 (microphone) and 0.754–3.928 (camera). Hennebert et al. [27] analysed 10 sensors on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1. A single-sensor analysis is presented, yielding min-entropy values of 0–7.85; motion sensors, e.g. accelerometer and vibration sensors, produced the highest entropy.
@@
Sensors have also been suggested as entropy sources in low-cost RNG designs for mobile devices. Suciu et al. [38] proposed combining GPS along with accelerometer, gyroscope and orientation sensors on an HTC Google Nexus One, passing NIST SP 800-22 tests [39], but without precise entropy estimates. Wallace et al. [40] explored a multi-sensor RNG design (accelerometer, gyroscope, microphone, WiFi, GPS, camera) with a non-standard entropy evaluation using 37 Android devices. Unlike dedicated TRNGs (e.g. ring oscillators, Johnson–Nyquist thermal noise, quantum sources [41]), mobile sensors depend heavily on user behaviour and environmental factors, introducing biases and correlations absent in controlled entropy sources. Existing sensor-based mechanisms largely overlook these dynamics, relying principally on heuristic or model evaluation metrics that do not account well for skewed distributions and other biases affecting predictability. In contrast, proposals of new entropy sources and authentication mechanisms commonly rely on worst-case metrics, such as min-entropy as stipulated in NIST SP 800-90B and AIS 20/31 [42, 43, 41, 24, 25]. Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
@@
Voris et al. [26] Acc. $H_{\infty}$ 3.1–11.4 True random number generation
Hennebert et al. [27] 10 (Various) $H_{\infty}$ 0–7.85 Harvesting entropy from physical mobile sensors
Lv et al. [28] Acc., Gyr., Mag. $H_{\infty}$ 0.593–5.876 Entropy source analysis on smartphone
Krhovják et al. [17] Cam., Mic. $H_{\{1,\infty\}}$ 0.5–5.376 Randomness sources in mobile devices
Li et al. [15] Mic., Light $H_{1}$ 32.3–38.5 Zero-interaction device pairing
Wu et al. [36] Mic., Light $H_{1}$ 51–54 Zero-interaction device pairing
This work 25 (Various) $H_{\{0,1,2,\infty\}}$ 8.1–23.9 ($H_{\infty}$) Systematic entropy analysis, including sensor combinatons, for security systems
@@ 2.2 Definitions
We draw attention to four special cases of $\alpha$ that are widely used in the literature. First, the Hartley (Max) Entropy (Eq. 2) is the logarithm of the number of possible outcomes with non-zero probability and serves as an upper bound.
@@
Second, Shannon entropy (Eq. 3) is the classical definition of entropy in information theory, corresponding to the limit of $H_{\alpha}$ as $\alpha\to 1$.
@@
We note that $H_{1}$ provides an average-case measure of uncertainty. It takes into account the entire distribution of outcomes; however, an adversary may only need to guess the most likely event to gain an advantage. Min-entropy (Eq. 5) is used as a conservative, worse-case metric, accounting for the least favorable distribution of outcomes; it is the value in the limit $\alpha\to\infty$:
@@
$H_{\alpha}$ is a non-increasing function of $\alpha$, i.e. $H_{\infty}(X)\leq\dots\leq H_{1}(X)\leq H_{0}(X)$. $H_{\infty}(X)$ focuses on the single most likely outcome, providing a strictly tighter (and generally minimum) bound on uncertainty. We observe that min-entropy ensures that even the most skewed probability distributions still meet the required security guarantees; indeed, it is a recommended method for assessing entropy sources in modern standards [24, 25]. We also rely on joint entropy for assessing the entropy of multiple random variables.
@@ 3.1 Threat Model and Assumptions
We consider an adversary aiming to compromise sensor-based security schemes (e.g., key generation, proximity detection, and continuous authentication). The adversary may gather extensive statistical data about how smartphone sensors behave in everyday usage; for instance, from widely available open datasets. We assume the attacker focuses on predicting or guessing the sensor outputs by prioritising the most probable values first, exploiting any biases in the distribution of sensor measurements. As such, they may resort to exhaustive enumeration of the measurement space if the target source’s entropy is low enough. We exclude capabilities such as fault injection and other hardware attacks (see [44]) as well as side-channel attacks such as cross-device correlation attacks [45]. While those could further reduce the effective entropy space—say, by inducing errors in the output values of sensing hardware—we regard them as outside the scope of this work. This threat model represents an adversary who can capitalise on statistical biases in sensor data without directly comprising the device physically. Our goal is to evaluate whether sensor data distributions offer sufficient entropy to resist attacks that search the space of sensor measurement values informed by their statistical properties.
@@ 3.2 High-level Methodology
In this work, we focus on estimating population-level (“global”) sensor distributions by pooling all users within a dataset. This choice reflects typical deployments in which systems must serve previously unseen users (e.g. first-time proximity checks, payment defence mechanisms, or ad-hoc device pairing) and thus cannot assume per-user enrollment or stable user-specific priors. Our approach involves five main stages:
@@
    Data acquisition. We collect large-scale sensor readings from public datasets spanning diverse activities and device usage. Modalities include motion, environmental, and orientation sensors (dataset details appear at the end of this section).
@@
    Interpretation. Finally, we interpret the resulting entropy measures, focusing on whether sensor outputs remain sufficiently unpredictable against an informed adversary. We compare single-sensor versus multi-sensor scenarios to verify if combining modalities truly alleviates biases or simply adds redundant data susceptible to predictability concerns.
Throughout, we note the constraints of NIST SP 800-90B, SP 800-22 [24], and AIS 20/31 [25] when applied to multi-sensor streams [37, 28]. These frameworks target univariate sources rather than joint entropy of complex multivariate processes, which is the focus of our work. (For example, NIST SP 800-90B focusses on assessing entropy sources with reduced, i.e. 8-bit, outputs [37, 24]). To scope suitable data, we searched IEEE DataPort, Google Scholar, Google Dataset Search, and GitHub. Several ostensibly “open” datasets were unavailable for download or had restrictive licensing [46, 47, 48]. We selected four datasets offering diverse contexts, consistent sampling, and documented modalities:
@@
    UCI-HAR [3]: A widely referenced dataset for human activity recognition, comprising smartphone sensor recordings from multiple subjects performing daily activities. Data includes triaxial accelerometer and gyroscope signals.
@@
    University of Sussex–Huawei Locomotion (SHL) [1, 2]: Sampled at 100 Hz from an Huawei Mate 9 smartphone. TWe use the publicly available SHL Preview dataset (three recording days per user; 59 hours total) and, for scope alignment with prior work, the handheld-phone subset.
@@
    Relay [5]: Approximately 1,500 NFC-based contactless transactions recorded at 100 Hz across several physical locations (e.g. cafés), including accelerometer, gyroscope, and environmental readings in emulated payment scenarios.
@@
    PerilZIS [49]: Collected at 10 Hz from a TI SensorTag, Samsung Galaxy S6, and Samsung Galaxy Gear, spanning multiple zero-interaction security use cases in an office environment.

### v10

@@ 1 Introduction
Modern mobile devices expose rich motion and ambient signals from accelerometers, gyroscopes, magnetometers, and more. Beyond activity recognition and context inference, e.g. gjoreski2018university , wang2019enabling , anguita2013public , yin2024systematic , a long line of work proposes using these signals for security-critical tasks, such as cryptographic key generation, zero-interaction device pairing, and continuous authentication gurulian2017effectiveness , gurulian2018good , mehrnezhad2015tap , mayrhofer2009shake , shrestha2014drone , shrestha2018sensor , shepherd2017applicability , shi2011senguard , miettinen2014conxsense , li2013unobservable , li2020t2pair , riva2012progressive , krhovjak2007sources , xu2021key , abuhamad2020sensor , mayrhofer2021adversary . Traditional key generation protocols mayrhofer2009shake , xu2021key have historically relied on shared motion patterns in order to establish secure communication between co-located devices. Meanwhile, other proposals have incorporated environmental phenomena, such as magnetic field characteristics and thermal fluctuations, to mitigate relay attacks shrestha2014drone , shrestha2018sensor , markantonakis2024using , gurulian2017effectiveness , shepherd2017applicability and to reduce the frequency of user authentication prompts krhovjak2007sources , riva2012progressive , shi2011senguard , miettinen2014conxsense , li2013unobservable .
@@
Despite these advances, a crucial assumption remains under-examined: that sensor data is robust against adversaries that aim to forge or reconstruct such data through statistical inference. Many existing proposals rely on heuristic assessments or model classification metrics, e.g. markantonakis2024using , mehrnezhad2015tap , gurulian2018good , shrestha2014drone , truong2014comparing , abuhamad2020sensor , mayrhofer2021adversary , which do not address two basic questions: (1) how much entropy do sensors actually provide and (2) to what extent do multimodal sensor combinations improve entropy? For the latter, multimodal sensing has been proposed as a means to counteract sensor-specific weaknesses shrestha2014drone , markantonakis2024using , truong2014comparing , mehrnezhad2015tap , but the quantitative security benefits of combining multiple sensors have not been rigorously evaluated. The security implications of low-entropy sensor data can be both direct and severe. If such data is used as authentication evidence or to derive secret keys, an adversary may attempt to enumerate the entire space of possible measurement values in order to (re-)derive it. Modern standards provide important quantitative thresholds for what can be considered ‘secure’ for such reasons. For example, NIST SP 800-57 nistsp80057 states that security levels below 80 bits are considered insecure for long-term use, with 112 bits as the current minimum and 128 bits recommended for cryptographic keys. Standards that govern random number generators (RNGs) are stricter still. NIST SP 800-90B turan2018recommendation effectively requires $\approx$1 bit of min-entropy per output bit ($\epsilon\leq 2^{-32}$), while AIS 20/31 bsi2024ais31 mandates over 240 and 250 bits of min-entropy and Shannon entropy respectively for an RNG’s internal state.
@@
In this paper, we analyse the entropy of 25 mobile sensors, which is substantially more than prior work—voris2011accelerometers (1 sensor), hennebert2013entropy (10), krhovjak2007sources (2), lv2020analysis (3)—across multiple public datasets. Our results reveal widespread weaknesses, both corroborating and expanding on existing results. Commodity sensors exhibit pronounced biases and dependencies, yielding on average 3.408–4.483 bits of min-entropy and 5.584–9.266 bits of Shannon entropy per sample across datasets. Combining multiple sensors confers only modest gains: we explore how inter-sensor correlations and statistical bias contribute to a divergence between average- and worst-case entropy, with min-entropy typically $\approx$40–75% below Shannon entropy for the same signals. Even when combining over 20 sensors simultaneously, the (joint) min-entropy remains below $\approx$22 bits, far below the thresholds considered secure by the aforementioned standards. In practical terms, schemes that treat these measurements as secrets expose a guess space small enough to be enumerable by an adversary. Our contributions are as follows:
@@ 2.1 Related Work
Mobile sensing has been widely explored for security tasks in three broad families: (1) proximity detection and relay-attack mitigation, where co-presence is inferred from shared motion or ambient context halevi2012secure , mehrnezhad2015tap , shepherd2017applicability , gurulian2017effectiveness , markantonakis2024using , gurulian2018good , shrestha2014drone ; (2) device pairing, where two devices derive evidence of co-location to establish a secure association or shared secret mayrhofer2009shake , pan2018universense , xu2021key , mayrhofer2021adversary ; and (3) continuous authentication, which aims to verify users passively during normal interaction riva2012progressive , shi2011senguard , miettinen2014conxsense , li2013unobservable , patel2016continuous , mekruksavanich2021deep , li2018sensor , micallef2015aren . Representative designs include tap-to-pair schemes using accelerometers mehrnezhad2015tap , induced shared vibrations gurulian2018good , and ambient-signal approaches that leverage temperature, humidity, magnetic field, or other environmental cues shrestha2014drone , markantonakis2024using . These mechanisms commonly fuse accelerometer and gyroscope measurements to capture correlated motion as evidence of co-presence mayrhofer2009shake , pan2018universense .
@@
Our key observation is that model-performance metrics are typically decoupled from the entropy of the underlying data sources upon which such models are based. Across the above application areas, effectiveness is usually inferred from discriminative performance, i.e. the system’s performance in classifying illegitimate vs. legitimate samples. A common pipeline—documented across proximity detection, pairing, and continuous authentication—collects data from $N$ participants; derives time- or frequency-domain features (e.g. cross-correlation, spectral energy, and Hamming, Euclidean, and mean absolute distances); learns a decision rule (thresholds or supervised models such as SVMs, Random Forests, Naïve Bayes); and reports FPR/FNR, EER, ROC/AUC, accuracy, or F1. This template appears in numerous systems, including mehrnezhad2015tap , shrestha2014drone , markantonakis2024using , shepherd2017applicability , riva2012progressive , li2013unobservable , miettinen2014conxsense , patel2016continuous , mekruksavanich2021deep , gurulian2018good . For a comprehensive treatment of the model evaluation metrics used by existing mobile sensor-based security systems, the reader is referred to the surveys by Abuhamad et al. abuhamad2020sensor , Alzubaidi and Kalita alzubaidi2016authentication , and Mayrhofer and Sigg mayrhofer2021adversary . Some proposals have attempted to assess the entropy of sensor signals within the context of a security mechanism, but this represents a small minority of the literature. Specifically, T2Pair by Li et al. li2020t2pair reports 32.3–38.5 bits of Shannon entropy, and a refinement by Wu et al. wu2024t2pair reports 51–54 bits. We are aware of no work that applies min-entropy as a worst-case measure of unpredictability turan2018recommendation , bsi2024ais31 , buller2016estimating , alzubaidi2016authentication , xu2021key , mayrhofer2021adversary . Adjacent evidence also highlights practical limits under tight timing constraints. For NFC transactions, EMV specifies $\approx$500 ms for proximity checks. In that setting, Markantonakis et al. markantonakis2024using (building on gurulian2017effectiveness , shepherd2017applicability ) report EERs of 0.179–0.246 across sensor combinations, concluding that commodity sensors are unsuitable for reliable proximity and relay detection without incurring usability or security issues. This points to wider fragility issues with sensor-based security schemes in practice.
@@
Some studies have attempted to assess the entropy bestowed by mobile sensors on a single-sensor basis. Voris et al. voris2011accelerometers investigated accelerometers as true random number generators (TRNGs) on a WISP RFID tag and Nokia N97 phone. The authors find that min-entropy is proportional to the motion applied to the device, with stationary movement having the lowest min-entropy. Intrinsic noise from the sensor circuitry and seismic activity, and the sampling rate of its analog-to-digital converter (ADC), are found to influence entropy generation. Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer. Lv et al. lv2020analysis analysed three mobile sensors—a triaxial accelerometer, gyroscope, and magnetometer—on an undisclosed Xiaomi Redmi smartphone. Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method. Krhovják et al. krhovjak2007sources examined image and audio data collected from mobile phone cameras and microphones, respectively, on a Nokia N73 and E-Ten X500 and M700 phones. The work reports Shannon entropies of 2.9 (microphone) and 2.408–5.376 (camera), and min-entropies of 0.5 (microphone) and 0.754–3.928 (camera). Hennebert et al. hennebert2013entropy analysed 10 sensors on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1. Min-entropy values of 0–7.85 are given, with motion sensors, e.g. accelerometer and vibration sensors, producing the highest entropy.
@@
Sensors have also been suggested as entropy sources in low-cost RNG designs for mobile devices. Suciu et al. suciu2011unpredictable proposed combining GPS along with accelerometer, gyroscope and orientation sensors on an HTC Google Nexus One, passing NIST SP 800-22 tests rukhin2001nist , but without precise entropy estimates. Wallace et al. wallace2016toward explored a multi-sensor RNG design (accelerometer, gyroscope, microphone, WiFi, GPS, camera) with a non-standard entropy evaluation using 37 Android devices. Unlike dedicated TRNGs (e.g. ring oscillators, Johnson–Nyquist thermal noise, quantum sources hurley2020quantum ), mobile sensors depend heavily on user behaviour and environmental factors, introducing biases and correlations absent in controlled entropy sources. Existing sensor-based mechanisms largely overlook these dynamics, relying principally on heuristic or model evaluation metrics that do not account well for skewed distributions and other biases affecting predictability. In contrast, proposals of new entropy sources and authentication mechanisms commonly rely on worst-case metrics, such as min-entropy as stipulated in NIST SP 800-90B and AIS 20/31 mai2017guessability , uellenbeck2013quantifying , hurley2020quantum , turan2018recommendation , bsi2024ais31 . Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
@@
Voris et al. voris2011accelerometers Acc. $H_{\infty}$ 3.1–11.4 True random number generation
Hennebert et al. hennebert2013entropy 10 (Various) $H_{\infty}$ 0–7.85 Harvesting entropy from physical mobile sensors
Lv et al. lv2020analysis Acc., Gyr., Mag. $H_{\infty}$ 0.593–5.876 Entropy source analysis on smartphone
Krhovják et al. krhovjak2007sources Cam., Mic. $H_{\{1,\infty\}}$ 0.5–5.376 Randomness sources in mobile devices
Li et al. li2020t2pair Mic., Light $H_{1}$ 32.3–38.5 Zero-interaction device pairing
Wu et al. wu2024t2pair Mic., Light $H_{1}$ 51–54 Zero-interaction device pairing
This work 25 (Various) $H_{\{0,1,2,\infty\}}$ 8.1–23.9 ($H_{\infty}$) Systematic entropy analysis, including sensor combinations, for mobile device sensors
@@ 2.2 Definitions
We use four cases of $\alpha$ that are widely used in the literature. First, the Hartley (Max) Entropy (Eq. 2) is the logarithm of the number of possible outcomes with non-zero probability and serves as an upper bound.
@@
Second, Shannon entropy (Eq. 3) is the classical definition of entropy in information theory, corresponding to the limit of $H_{\alpha}$ at $\alpha\to 1$.
@@
We note that $H_{1}$ provides an average-case measure of uncertainty. It takes into account the entire distribution of outcomes; however, an adversary may only need to guess the most likely event to gain an advantage. To account for this, min-entropy (Eq. 5) is used as a conservative, worse-case metric, accounting for the least favorable distribution of outcomes; it is the value in the limit $\alpha\to\infty$:
@@
$H_{\alpha}$ is a non-increasing function of $\alpha$, i.e. $H_{\infty}(X)\leq\dots\leq H_{1}(X)\leq H_{0}(X)$. $H_{\infty}(X)$ focuses on the single most likely outcome, providing a strictly tighter (and generally minimum) bound on uncertainty. We observe that min-entropy ensures that even the most skewed probability distributions still meet the required security guarantees; indeed, it is a recommended method for assessing entropy sources in modern standards turan2018recommendation , bsi2024ais31 . We also rely on joint entropy for assessing the entropy of multiple random variables.
@@ 3.1 Threat Model and Assumptions
We consider an adversary aiming to compromise sensor-based security schemes (e.g. key generation, proximity detection, and continuous authentication). The adversary may gather extensive statistical data about how smartphone sensors behave in everyday usage; for instance, from widely available open datasets. We assume the attacker focuses on predicting or guessing the sensor outputs by prioritising the most probable values first, exploiting any biases in the distribution of sensor measurements. As such, they may resort to exhaustive enumeration of the measurement space if the target source’s entropy is low enough. We exclude capabilities such as fault injection and other hardware attacks (see shepherd2021physical ) as well as side-channel attacks such as cross-device correlation attacks dautov2019effects . While those could further reduce the effective entropy space—say, by inducing errors in the output values of sensing hardware—we regard them as outside the scope of this work. This threat model represents an adversary who can capitalise on statistical biases in sensor data without directly comprising the device physically. Our goal is to evaluate whether sensor data distributions offer sufficient entropy to resist attacks that search the space of sensor measurement values informed by their statistical properties.
@@ 3.2 High-level Methodology
In this study, we focus on estimating population-level (“global”) sensor distributions by pooling all users within a dataset. This choice reflects typical deployments in which systems must serve previously unseen users (e.g. first-time proximity checks, payment defence mechanisms, or ad-hoc device pairing) and thus cannot assume per-user enrollment or stable user-specific priors. Our approach involves five main stages:
@@
    Data acquisition. We collect large-scale sensor readings from public datasets spanning diverse activities and device usage. Modalities include motion, environmental, and orientation sensors (dataset details are given at the end of this section).
@@
    Evaluation. Finally, we interpret the resulting entropy measures, focusing on whether sensor outputs remain sufficiently unpredictable against an informed adversary. We compare single-sensor versus multi-sensor scenarios to verify if combining modalities truly alleviates biases or simply adds redundant data susceptible to predictability concerns.
Throughout, we note the constraints of NIST SP 800-90B, SP 800-22 turan2018recommendation , and AIS 20/31 bsi2024ais31 when applied to multi-sensor streams buller2016estimating , lv2020analysis . These frameworks target univariate sources rather than joint entropy of complex multivariate processes, which is the focus of our work. (For example, NIST SP 800-90B focusses on assessing entropy sources with reduced, i.e. 8-bit, outputs buller2016estimating , turan2018recommendation ). To scope suitable data, we searched IEEE DataPort, Google Scholar, Google Dataset Search, and GitHub. Several ostensibly “open” datasets were unavailable for download or had restrictive licensing Mahbub_Btas2016_UMDAA02 , stragapede2023behavepassdb , acien2021becaptcha . We selected four datasets offering diverse contexts, consistent sampling, and documented modalities:
@@
    UCI-HAR anguita2013public : A widely referenced dataset for human activity recognition, comprising smartphone sensor recordings from multiple subjects performing daily activities. Data includes triaxial accelerometer and gyroscope signals.
@@
    University of Sussex–Huawei Locomotion (SHL) gjoreski2018university , wang2019enabling : Sampled at 100 Hz from an Huawei Mate 9 smartphone. We use the publicly available SHL Preview dataset containing three recording days per user (59 hours total). For scope alignment with prior work, the handheld-phone subset is used.
@@
    Relay gurulian2017effectiveness : Approximately 1,500 NFC-based contactless transactions recorded at 100 Hz across several physical locations (e.g. cafés), including accelerometer, gyroscope, and environmental readings taken from emulated payment scenarios.
@@
    PerilZIS fomichev2019perils : Measurements collected at 10 Hz from a TI SensorTag, Samsung Galaxy S6, and Samsung Galaxy Gear, spanning multiple zero-interaction security use cases in an office environment.